OpenStudy (anonymous):

Evaluate the expression 3.14(a2 + ab) when a = 4 and b = 2.

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

3.14(8+8) 3.14(16) =______

OpenStudy (anonymous):

50.24

OpenStudy (misssunshinexxoxo):

Solve 3.14×(42+4×3)

OpenStudy (misssunshinexxoxo):

http://www.wolframalpha.com/input/?i=3.14%284%5E2%2B4*3%29

OpenStudy (misssunshinexxoxo):

Answer is in link ^

OpenStudy (misssunshinexxoxo):

87.92

OpenStudy (anonymous):

wait did you mean a^2 or ax2

OpenStudy (misssunshinexxoxo):

a^2

OpenStudy (anonymous):

then outdent it be 3.14(16+8) since a= 4 and 4^2 = 16..and ab = 4(2) = 8

OpenStudy (anonymous):

wouldn't*

OpenStudy (anonymous):

is it 50.24 or 87.92?

OpenStudy (anonymous):

it not 50.24... it's 75.36

OpenStudy (anonymous):

i was wrong the first time since i thought you meant ax2 not a squared lol

OpenStudy (anonymous):

it is a x 2

OpenStudy (anonymous):

@haiimawesome

OpenStudy (anonymous):

its a to the power of 2

OpenStudy (anonymous):

so then that would be 3.14( 4^2 + 4(2)) = 3.14 (16 + 8) = 3.14(24) =_____

OpenStudy (anonymous):

75.36

OpenStudy (anonymous):

yep!

OpenStudy (anonymous):

thanks! c:

OpenStudy (anonymous):

np :)